OpenStudy (anonymous):

Solve K=1/3 (b-2a) for b. I dont even know where to start... Sorry for the continuation of questions.

OpenStudy (katrinakaif):

3k + 2a = b

OpenStudy (katrinakaif):

Bring the 1/3 to the other side by multiplying the reciprocal of it with "k" so reciprocal of 1/3 = 3 and multiply by k = k x 3 = 3k Now your left with 3k = b - 2a Bring the 2a to the other side by adding it.. 3k + 2a = b

OpenStudy (nilankshi):

good kaif

OpenStudy (katrinakaif):

Thanks =)

OpenStudy (anonymous):

\[K=b-2a/3\] \[3K=b-2a\] \[3K+2a=b\]

OpenStudy (nilankshi):

good kush

OpenStudy (anonymous):

got it MrBonez321

OpenStudy (anonymous):

thanks nilankshi di , well please come to chat box

OpenStudy (nilankshi):

ok